# DROP: Poison Dilution via Knowledge Distillation for Federated Learning

### v1

@@ 5.2. Results
5.2.1. Robustness Across FL Learning Configurations
A key strength of DROP is its robustness across diverse FL learning configurations, ensuring reliable protection against targeted backdoor attacks independent of the specific learning parameter settings. While other defenses exhibit significant performance variability depending on the enforced learning parameters (Figure 3), DROP consistently maintains near-zero ASR across all configurations of interest, as shown in Table 4. This consistency demonstrates DROP’s adaptability and makes it a learning configuration agnostic defense, capable of handling varying client learning rates, batch sizes, and local training epochs without compromising on security. This robustness is particularly important for real-world FL deployments, where learning configurations can vary dynamically across tasks. By ensuring reliable performance across a wide range of setups, DROP offers practical, consistent protection against adversarial threats, making it well-suited for heterogeneous FL environments.
@@ 5.2.1. Robustness Across FL Learning Configurations
Table 4. MTA (%) and ASR (%) for DROP on CIFAR-10 for 1.25% DPR and 20% MCR, across all 10 configurations identified in Section 3.2. DROP achieves low ASR across all configurations.
Figure 5. MTA (a) and ASR (b) across rounds for various defenses, for CIFAR-10 with 1.25% DPR and 20% MCR for configuration C4. Certain defenses like FLIP and FLAME have wildly fluctuating ASR across rounds, making them unreliable. DROP instead achieves consistently low ASR in all rounds.
Consistency Across Rounds. A robust defense must ensure low ASR during rounds where the MTA remains acceptable, preventing randomness in ASR fluctuations from determining the outcome of FL training. While some baselines (e.g., FLIP, FLAME) initially achieve low ASR, their performance deteriorates over subsequent rounds, resulting in inconsistent suppression of the attack. This variability is problematic because in a real-world setting, FL training might terminate at any round, and reliance on transient phases of low ASR does not guarantee reliable protection. In Figure 5, we compare the performance of various defensive baselines with DROP over 200 rounds. Although some baselines degrade more quickly (e.g., Median, FLTrust), others (e.g., FLIP, FLAME) exhibit slower degradation, yet suffer from significant ASR fluctuations across rounds. This inconsistency makes it difficult to guarantee robust defense throughout the training process. In contrast, DROP consistently maintains near-zero ASR across all rounds, offering reliable protection regardless of when training might terminate. As described in the Property 2 (Section 2.2), this stability ensures that practitioners do not have to depend on chance for effective defense, making DROP a practical and dependable solution for real-world FL systems.
5.2.2. Adaptability to Varying Attack Stealthiness
@@
Varying DPR. By varying the attacker’s DPR, we simulate increasingly stealthy targeted backdoor attacks, with DPR as low as 1.25%. We observe that all baseline defenses fail to defend against targeted backdoors across CIFAR-10, CINIC-10, and EMNIST, regardless of attack stealthiness (Table 3). Some defenses perform adequately only when the attack is overt. For instance, for CIFAR-10, FLARE mitigates the attack effectively at 5% DPR, achieving 3.5% ASR, but fails entirely when the DPR is reduced (e.g., 2.5%). Similarly, for CINIC-10, FLIP demonstrates moderate robustness at 2.5% DPR, achieving 38% ASR, but succumbs to the attack at 1.25% DPR. In contrast, DROP consistently defends against targeted backdoor attacks across all levels of DPR, maintaining near-zero ASR (1.5%, 1.2%, and 0.5%) in all configurations.
@@
MCR (%) 20 40 20 40 20 40 20 40 20 40 20 40
Undefended 86.56 86.05 92.1 99.6 74.49 74.47 98.65 99.6 87.78 87.65 94.75 100.0
FLAME 80.78 81.11 91.3 99.8 72.77 72.62 98.6 99.32 86.43 86.71 2.50 89.50
FLARE 83.28 83.74 58.3 98.9 71.09 71.65 93.02 97.27 86.72 86.52 27.50 96.50
FLIP 82.25 83.43 11.2 72.3 69.6 71.65 38.82 92.94 87.48 87.54 12.25 84.75
FLTrust 87.38 87.24 99.5 99.4 74.54 74.65 98.72 99.73 88.47 88.68 78.75 99.50
Fool’s Gold 83.80 83.73 86.2 99.7 69.96 69.93 95.29 98.77 85.91 85.82 25.50 96.75
Median 86.10 85.54 98.9 99.1 73.38 73.26 98.87 99.69 87.66 87.71 94.25 99.75
Multi-KRUM 84.41 84.04 99.2 99.8 73.58 73.58 99.74 99.55 87.13 87.20 1.00 99.25
DROPlet (ours) 85.49 84.87 0.9 0.8 74.27 74.65 0.48 99.73 87.66 87.37 1.50 0.75
DROP (ours) 69.11 70.71 1.3 1.4 65.72 64.97 1.22 2.36 88.67 88.43 0.75 0.75
@@
Table 5. MTA (%) and ASR (%) metrics at different Malicious Client Ratios (MCR) for CIFAR-10, CINIC-10 and EMNIST datasets. The DPR is fixed at 2.5%. Existing defenses fail to prevent the attack, particularly at 40% MCR. DROP is resilient in all cases, despite some reduction in MTA. DROPlet performs adequately in most cases independently of the MCR. The learning configurations are C1 for CIFAR-10 (C4 for CINIC-10) and C9 for EMNIST. Refer to Table 2 and Table 7 for learning configuration details.
@@
Varying MCR. By varying the MCR, we simulate scenarios where the attacker controls a smaller (20%) or larger (40%) fraction of clients in the federation. A high MCR, such as 40%, significantly increases the likelihood that malicious clients will outnumber benign ones in certain FL rounds due to the random client selection process. This imbalance enables adversarial updates to dominate the aggregation process, amplifying the attack’s effectiveness and exposing vulnerabilities in clustering-based defenses. Once again, we observe that most defensive baselines fail to provide adequate protection against the attack under such conditions (Table 5). For instance, FLIP reduces the ASR to 11.2% and 38.82% for 20% MCR on CIFAR-10 and CINIC-10, but fails to mitigate the attack at 40% MCR. FLAME, due to its overly aggressive adaptive clipping scheme, struggles to achieve convergence. Both FLAME and Multi-Krum achieve resilience for 20% MCR on EMNIST, but do not thwart the attack at 40% MCR. In contrast, DROP consistently demonstrates resilience across both 20% and 40% MCR setups, achieving near-zero ASR while, highlighting its robustness even in highly adversarial settings.
5.2.3. Robustness - Accuracy Tradeoff
@@
While DROP demonstrates strong effectiveness against backdoor attacks, it comes with a notable trade-off in MTA compared to some baselines. For example, in Table 3 with a DPR of 1.25%, DROP achieves an MTA of 74.74%, which is lower than FLTrust (87.93%) and Median (87.14%). Similarly, in Table 5 with an MCR of 20%, DROP records an MTA of 68.74%, compared to 86.88% for FLTrust. This reduction in accuracy primarily stems from the strict distillation-based cleansing mechanism, which aims to ensure near-total suppression of adversarial influences. Notably, unlike clustering or aggregation-only defenses, DROP applies knowledge distillation by using a model stealing technique (i.e., MAZE), which contributes significantly to the observed MTA drop while exhibiting strong robustness. The trade-off between robustness and accuracy is a well-documented challenge in adversarial machine learning (Tsipras et al., 2019).
@@
DROPlet - Lightweight Variant of DROP. DROPlet (Section 4.4) is designed for environments where computational efficiency is a priority. Unlike DROP, DROPlet omits the knowledge distillation component, thereby eliminating the associated computational overhead. This makes it particularly suitable for FL deployments with limited resources or scenarios where minimizing server-side processing time is critical. Notably, execution time for a single non-IID run of DROPlet is comparable to other baselines, completing in approximately 30 minutes, whereas DROP takes 2 hours, and FLIP requires an extensive 57 hours for the same setup.
While DROPlet may not perform as well as DROP against certain highly stealthy, low-DPR attacks (e.g., CIFAR-10, 1.25% DPR in Table 3) or high-MCR attacks (e.g., CINIC-10, 40% MCR in Table 5), it still provides strong protection in many practical configurations, achieving near-zero ASR in less covert attack scenarios and maintaining robustness across a broad range of learning setups (Table 3 and Table 5). For simpler or moderately stealthy attack scenarios with substantial adversarial presence within the federation, DROPlet performs comparably to DROP without the additional computational cost incurred by the knowledge distillation step.
However, it is important to note that an adaptive adversary aware of this defense could lower its DPR to evade detection, exploiting the absence of the distillation component to bypass DROPlet’s defenses. This limitation makes DROPlet less suitable for environments where adversaries are likely to adopt highly stealthy strategies. Nevertheless, its lightweight design offers significant practical value as a plugin to existing FL defenses, requiring minimal code modifications while enhancing robustness without incurring substantial overhead. By striking a balance between computational efficiency and robustness, DROPlet provides a viable defense solution for resource-constrained FL settings.
@@
Table 6. MTA and ASR across rounds for various defenses for non-IID data with varying $Dir(\alpha)$, for 1.25% DPR and 20% MCR, using configuration C4 for CIFAR-10 and C9 for EMNIST. FLAME fluctuates wildly and often ends up with models that have not learned anything effectively, leading to random-guess MTA (red) and consequently, zero ASR. Results for FLIP are shown only for $Dir(\alpha)=1$ due to its extreme computational overhead, as the method requires adversarial training for each client at every round. As observed, DROP is the most resilient defense on EMNIST and all defenses struggle on the more complex CIFAR-10 learning task.
5.2.4. Hardness Under Non-IID Settings
@@
Table 6 presents the results of evaluating various defenses under increasingly non-IID conditions for CIFAR-10 and EMNIST. While both datasets are commonly used in FL research, CIFAR-10 poses a greater challenge due to its higher intra-class diversity, which allows malicious updates to blend more seamlessly with legitimate ones, making it harder to detect and defend against backdoor attacks. This complexity is reflected in the relatively poor performance of most defenses on CIFAR-10 under non-IID conditions. FLIP for instance, while reducing ASR compared to most other defenses, still suffers from significant fluctuations in ASR, sometimes exceeding 60% under non-IID conditions. This variability highlights the challenge of achieving reliable defense in high-stakes applications where consistency is crucial.
@@
While our defense performs well in many non-IID scenarios, the fluctuating ASR observed in more complex tasks like CIFAR-10 underscores the difficulty of defending against backdoor attacks under such conditions. These results suggest that more research is needed to develop defenses capable of better identifying malicious updates in the presence of significant data heterogeneity.
6. Conclusion
@@
To address this gap, we propose DROP, a novel defense mechanism that is agnostic to learning configurations and the adversary’s stealthiness. By integrating clustering, activity tracking, and knowledge distillation, DROP offers a comprehensive solution for mitigating targeted backdoor attacks, ensuring reliable protection across diverse FL setups. Our extensive evaluation across multiple datasets, varying data distributions, and a wide range of attacker configurations highlights DROP’s robustness. Unlike existing baselines, which often fail under certain learning configurations or in the presence of stealthy attacks, DROP consistently achieves near-zero attack success rate despite a drop in main task accuracy.
Limitations. Although DROP’s additional compute overhead from the knowledge distillation step is manageable, it is still significant compared to lightweight baselines. However, this overhead is notably lower than client-side approaches like FLIP, which involve adversarial training. Second, despite the near-total mitigation of backdoor attacks, the reduction in MTA may not be acceptable in certain high-utility applications. However, it is worth noting that DROPlet, our more lightweight variant, offers a balance between robustness and accuracy, while being more computationally efficient than DROP. Our results also reveal that defending against targeted backdoor attacks in complex learning tasks like CIFAR-10, particularly in highly non-IID settings, remains an open challenge. This highlights the need for further research into adaptive, configuration-independent defenses capable of handling evolving adversarial strategies.
@@ Acknowledgments
Appendix A Bound Analysis on Number of Malicious Clients per Round
Let $N$ denote the total number of clients and $M$ the number of malicious clients. The ratio of malicious clients is given by the Malicious Client Ratio (MCR), $\rho=\frac{M}{N}$. In each round of Federated Learning (FL), we sample $C$ clients randomly from the $N$ total clients. Let $\mathcal{M}$ be the number of malicious clients (out of $C$) sampled in some round. We are interested in finding a lower bound on the probability that there are more malicious clients than benign ones in a round of FL, i.e., $P\left(\mathcal{M}\geq\frac{C}{2}\right)$.
The selection of malicious clients can be modeled as a Binomial distribution where probability of success is the MCR ($\rho$). Thus, our desired probability can be written in terms of the CDF $F$ of this distribution:
(11) $P\left(\mathcal{M}\geq\frac{C}{2}\right)=1-F\left(\frac{C}{2},C,\rho\right).$
Using a Chernoff bound (Arratia and Gordon, 1989), we can thus obtain a lower bound on the probability that malicious clients outnumber benign ones in a given round as:
(12) $P\left(\mathcal{M}\geq\frac{C}{2}\right)$ $\geq 1-\exp\left(-C\cdot\left(\frac{1}{2}\left(\ln\frac{1}{2\rho}+\ln\frac{1}{2(1-\rho)}\right)\right)\right)$
(13) $=1-\exp\left(C\cdot\frac{1}{2}\ln\left(4\rho(1-\rho)\right)\right)$
(14) $=1-\left(4\rho(1-\rho)\right)^{\frac{C}{2}}.$
Appendix B Related Work
The increasing adoption of FL has spurred extensive research on defending against poisoning attacks, particularly backdoor attacks, which pose a serious threat due to their ability to embed hidden malicious behaviors without degrading the model’s overall performance. Several defense mechanisms have been proposed to mitigate these threats, including robust aggregation methods, trust-based filtering, and anomaly detection techniques. Despite these efforts, existing defenses often exhibit limitations when evaluated across diverse attack configurations, especially under varying data poisoning rates, malicious client ratios within the federation and non-IID client data distributions.
Robust aggregation-based defenses. Coordinate-wise median aggregation (Yin et al., 2018) and Multi-Krum (Blanchard et al., 2017) are two classical approaches designed to mitigate Byzantine failures by robustly aggregating model updates. The coordinate-wise median computes the median value for each model parameter across client updates, neutralizing the impact of outliers. Multi-Krum, on the other hand, iteratively selects and aggregates updates that are closest to the majority based on pairwise distances, making it resilient to adversarial updates. While these methods are effective against simple poisoning attacks, they struggle to defend against stealthy backdoor attacks, particularly when the data poisoning rate is low or the client data distribution is highly skewed. Additionally, both methods assume IID client data, limiting their robustness in realistic non-IID FL settings.
Trust-based defenses. FLTrust (Cao et al., 2021) introduces a server-side reference model trained on a small trusted dataset to measure the trustworthiness of client updates. Only updates that align closely with the reference model are aggregated, providing a strong baseline against various adversarial strategies. However, FLTrust’s performance can degrade when the trusted dataset is not fully representative of the overall data distribution and if the attack is very stealthy.
Similarity and anomaly detection-based defenses. Fool’s Gold (Fung et al., 2018) uses similarity-based clustering to detect and penalize clients that contribute updates with similar gradients across multiple rounds, under the assumption that adversaries tend to behave similarly. While this approach reduces the contribution of malicious clients, its reliance on similarity metrics makes it susceptible to adaptive adversaries that can evade detection by introducing slight randomness in their updates. FLAME (Nguyen et al., 2022), which employs adaptive clipping to detect and filter anomalous gradients, also shows promise in defending against targeted attacks. However, as we observe in our experiments, FLAME’s adaptive clipping can become overly aggressive, preventing convergence in some scenarios.
Penultimate layer representation-based defenses. FLARE (Wang et al., 2022) leverages discrepancies in penultimate layer representations (PLR) of model updates to assign trust scores and filter out potentially malicious updates. This approach works well for overt poisoning attempts but, as we demonstrate, struggles to mitigate stealthy attacks with low DPR, where poisoned updates closely resemble benign ones.
Adversarial training-based and sample rejection defenses. FLIP (Zhang et al., 2023) represents a different class of defenses by introducing the ability to reject individual samples during aggregation. By leveraging adversarial training and low-confidence refusals, FLIP aims to reconstruct client-side triggers and reject poisoned updates. While effective to some extent, FLIP’s rejection mechanism often discards a non-negligible fraction of clean samples, which can negatively impact the main task accuracy (MTA). As observed in our experiments, this trade-off in rejection-based defenses can lead to deceptively high MTA metrics, as benign samples that are incorrectly rejected are excluded from evaluation. In contrast, our proposed defense does not require such sample-level rejection, thereby preserving the overall integrity of the clean data and achieving a better balance between MTA and ASR across varying configurations. A crucial consideration when designing FL frameworks is that participating nodes often have limited computational resources, which can constrain the feasibility of complex operations. e.g., FLIP relies on client-side adversarial training—a computationally expensive process. While this approach is viable in their setup, where only 10 clients are sampled per round, increasing the number of selected clients significantly extends the duration of each round, making it impractical for large-scale deployments.
Despite the progress made by these defenses, a significant gap remains: no existing method consistently performs well across different DPR levels, IID and non-IID data distributions, and varying MCR. Our proposed defense, DROP, addresses these issues by providing a robust, configuration-agnostic solution that is effective across a broad range of attack and learning configurations.
Appendix C Experimental Details
C.1. Baselines
Each of the defense methods which were presented aims to limit the influence of malicious updates during aggregation. However, most works provide limited or inconsistent details about their evaluation setups, particularly concerning client learning configurations such as learning rate, batch size, and the number of local training epochs. For instance, Median (Yin et al., 2018) is evaluated on a simpler learning task (MNIST (Deng, 2012)) and specifies only the total number of participating clients, without providing key details about the client learning setup, such as the learning rate, batch size, or the number of local epochs. Similarly, Multi-Krum (Blanchard et al., 2017) reduces the impact of outliers using robust statistics but primarily focuses on how the data is partitioned among clients. While it does evaluate the method across different batch sizes, it lacks a detailed discussion of the broader local training setup. FLTrust (Cao et al., 2021) adopts a trusted server-side reference model to filter anomalous updates. The authors report using a ”combined” learning rate of 0.002, a batch size of 64, and a single local training epoch. In contrast, FoolsGold (Fung et al., 2018), which identifies and penalizes suspiciously similar client contributions, does not explicitly report the learning rate or the number of local epochs in its evaluation, only mentioning batch sizes of 10 or 50 depending on the dataset. FLAME (Nguyen et al., 2022), which employs anomaly detection to flag potentially malicious gradients, describes the structure of the federation but omits critical information about the local learning setup, such as learning rate, batch size, or the number of epochs. FLIP (Zhang et al., 2023), on the other hand, presents a more detailed setup. We adapted the original implementation and hyperparameters for our evaluation. The original defense assumes that both the defense and malicious client poisoning are triggered after model convergence, citing interference in convergence if malicious activity begins earlier. However, our observations show that targeted backdoors do not disrupt model convergence even when initiated at the beginning of FL training. Thus, starting the defense only after convergence is ineffective since poisoning has already occurred by that point. On the other hand, initiating the defense too early leads to suboptimal performance. Therefore, we activate the defense after the first 10 rounds for CIFAR-10 and after the first 3 rounds for EMNIST to balance effectiveness and performance.
C.2. DROP Parameters
Knowledge distillation, particularly in the context of model stealing attacks, is inherently imperfect and cannot replicate the target model exactly, resulting in a minor decrease in MTA. To address this and ensure convergence in the FL setting, the knowledge distillation component is applied every $K$ rounds instead of every round, allowing the system to recover lost MTA during intermediate rounds. The budget parameter in model stealing attacks and knowledge distillation determines the number of queries used to generate synthetic samples, which are then employed to guide the distillation process. A sufficient query budget ensures the generation of high-quality synthetic data that aligns closely with the target model’s decision boundaries, thereby enhancing the effectiveness of knowledge distillation. For CIFAR-10, we set $K=5$ with a query budget of 5M queries. For EMNIST, $K=40$ with a query budget of 4M queries. These values strike a balance between computational efficiency and the quality of the distilled global model.
C.3. EMNIST Grid-Search
Figure 6. Visualizing the impact of the FL setup (particularly the learning-rate, batch-size, and number of epochs used by clients) on main-task accuracy attack success rate when poisoned clients aim to inject a targeted backdoor for the EMNIST dataset.
In the same fashion as Section 3.1 for CIFAR-10, we conduct a grid-search analysis over key hyperparameters, varying the client’s learning rate, batch size, and number of epochs on the EMNIST (Cohen et al., 2017) dataset. Our findings in Figure 6 indicate that targeted backdoor attacks are more likely to succeed across a wider range of learning parameter combinations, with numerous setups yielding an ASR greater than 80%. In Table 7, we highlight ten specific learning configurations where the attack achieves high ASR while maintaining a high MTA, underscoring the vulnerability of these setups to adversarial manipulation.
Config LR BS Epochs MTA (%) ASR (%)
C1 0.1 32 2 89.23 99.00
C2 0.1 64 5 88.22 99.00
C3 0.05 32 5 88.48 98.75
C4 0.1 32 1 89.59 98.75
C5 0.1 64 1 89.20 98.50
C6 0.1 32 5 88.73 98.50
C7 0.1 64 2 88.87 98.25
C8 0.05 32 1 89.21 98.25
C9 0.01 32 2 87.73 97.25
C10 0.025 128 5 86.61 96.00
Table 7. Client FL configurations for successful stealthy attacks on EMNIST i.e., cases where the MTA $\geq 80\%$ and ASR $\geq 95\%$.

### v2

@@ V Experiments
V-A Robustness Across FL Learning Configurations
A key strength of DROP is its robustness across diverse FL learning configurations, ensuring reliable protection against targeted backdoor attacks independent of the specific learning parameter settings. While other defenses exhibit significant performance variability depending on the enforced learning parameters (Figure 3), DROP consistently maintains near-zero ASR across all configurations of interest, as shown in Table V. This consistency demonstrates DROP’s adaptability and makes it a learning configuration agnostic defense, capable of handling varying client learning rates, batch sizes, and local training epochs without compromising on security. This robustness is particularly important for real-world FL deployments, where learning configurations can vary dynamically across tasks. By ensuring reliable performance across a wide range of setups, DROP offers practical, consistent protection against adversarial threats, making it well-suited for heterogeneous FL environments.
@@ V-A Robustness Across FL Learning Configurations
TABLE V: MTA (%) and ASR (%) for DROP on CIFAR-10 for 1.25% DPR and 20% MCR, across all 10 configurations identified in Section III-B. DROP achieves low ASR across all configurations.
Figure 5: MTA (a) and ASR (b) across rounds for various defenses, for CIFAR-10 with 1.25% DPR and 20% MCR for configuration C4. Certain defenses like FLIP and FLAME have wildly fluctuating ASR across rounds, making them unreliable. DROP instead achieves consistently low ASR in all rounds.
Consistency Across Rounds. A robust defense must ensure low ASR during rounds where the MTA remains acceptable, preventing randomness in ASR fluctuations from determining the outcome of FL training. While some baselines (e.g., FLIP, FLAME) initially achieve low ASR, their performance deteriorates over subsequent rounds, resulting in inconsistent suppression of the attack. This variability is problematic because in a real-world setting, FL training might terminate at any round, and reliance on transient phases of low ASR does not guarantee reliable protection. In Figure 5, we compare the performance of various defensive baselines with DROP over 200 rounds. Although some baselines degrade more quickly (e.g., Median, FLTrust), others (e.g., FLIP, FLAME) exhibit slower degradation, yet suffer from significant ASR fluctuations across rounds. This inconsistency makes it difficult to guarantee robust defense throughout the training process. In contrast, DROP consistently maintains near-zero ASR across all rounds, offering reliable protection regardless of when training might terminate. As described in the Property 2 (Section II-B), this stability ensures that practitioners do not have to depend on chance for effective defense, making DROP a practical and dependable solution for real-world FL systems.
V-B Adaptability to Varying Attack Stealthiness
@@
Varying DPR. By varying the attacker’s DPR, we simulate increasingly stealthy targeted backdoor attacks, with DPR as low as 1.25%. We observe that all baseline defenses fail to defend against targeted backdoors across CIFAR-10, CINIC-10, and EMNIST, regardless of attack stealthiness (Table III). Some defenses perform adequately only when the attack is overt. For instance, for CIFAR-10, FLARE mitigates the attack effectively at 5% DPR, achieving 3.5% ASR, but fails entirely when the DPR is reduced (e.g., 2.5%). Similarly, for CINIC-10, FLIP demonstrates moderate robustness at 2.5% DPR, achieving 38% ASR, but succumbs to the attack at 1.25% DPR. In contrast, DROP consistently defends against targeted backdoor attacks across all levels of DPR, maintaining near-zero ASR (1.5%, 1.2%, and 0.5%) in all configurations.
@@
MCR (%) NA 20 40 20 40 NA 20 40 20 40 NA 20 40 20 40
Undefended 86.48 86.56 86.05 92.1 99.6 74.71 74.49 74.47 98.65 99.6 87.74 87.78 87.65 94.75 100.0
FLAME 80.66 80.78 81.11 91.3 99.8 72.93 72.77 72.62 98.6 99.32 86.28 86.43 86.71 2.50 89.50
FLARE 83.63 83.28 83.74 58.3 98.9 71.07 71.09 71.65 93.02 97.27 87.05 86.72 86.52 27.50 96.50
FLIP 80.09 82.25 83.43 11.2 72.3 67.73 69.6 71.65 38.82 92.94 87.24 87.48 87.54 12.25 84.75
FLTrust 86.85 87.38 87.24 99.5 99.4 74.46 74.54 74.65 98.72 99.73 88.59 88.47 88.68 78.75 99.50
Fool’s Gold 83.93 83.80 83.73 86.2 99.7 70.25 69.96 69.93 95.29 98.77 85.96 85.91 85.82 25.50 96.75
Median 86.41 86.10 85.54 98.9 99.1 73.29 73.38 73.26 98.87 99.69 87.66 87.66 87.71 94.25 99.75
Multi-KRUM 84.28 84.41 84.04 99.2 99.8 73.61 73.58 73.58 99.74 99.55 87.37 87.13 87.20 1.00 99.25
DROPlet (ours) 85.69 85.49 84.87 0.9 0.8 74.62 74.27 74.65 0.48 99.73 87.77 87.66 87.37 1.50 0.75
DROP (ours) 69.7 69.11 70.71 1.3 1.4 65.4 65.72 64.97 1.22 2.36 88.77 88.67 88.43 0.75 0.75
@@
TABLE VI: MTA (%) and ASR (%) metrics under no attack (NA) and at different Malicious Client Ratios (MCR) for CIFAR-10, CINIC-10 and EMNIST datasets. The DPR is fixed at 2.5%. Existing defenses fail to prevent the attack, particularly at 40% MCR. DROP is resilient in all cases, despite some reduction in MTA. DROPlet performs adequately in most cases independently of the MCR. The learning configurations are C1 for CIFAR-10 (C4 for CINIC-10) and C9 for EMNIST. Refer to Table II and Table VIII for learning configuration details.
@@
Varying MCR. By varying the MCR, we simulate scenarios where the attacker controls a small (20%) or large (40%) fraction of clients. A high MCR like 40% significantly increases the likelihood that malicious clients will outnumber benign ones in certain FL rounds due to the random client selection process. This imbalance enables adversarial updates to dominate the aggregation process, amplifying the attack’s effectiveness and exposing vulnerabilities in clustering-based defenses. Once again, we observe that most defensive baselines fail to provide adequate protection against the attack under such conditions (Table VI). For instance, FLIP reduces the ASR to 11.2% and 38.82% for 20% MCR on CIFAR-10 and CINIC-10, but fails to mitigate the attack at 40% MCR. FLAME, due to its overly aggressive adaptive clipping scheme, struggles to achieve convergence. Both FLAME and Multi-Krum achieve resilience for 20% MCR on EMNIST, but do not thwart the attack at 40% MCR. In contrast, DROP consistently demonstrates resilience across both 20% and 40% MCR setups, achieving near-zero ASR while, highlighting its robustness even in highly adversarial settings.
Varying Attacks. DROP is resilient to stealthy state-of-the-art backdoor attack strategies like Neurotoxin and Chameleon. As shown in Table IV, it consistently maintains near-zero ASR (1.87% or lower) across all datasets and poisoning rates, even when facing these durable and adaptive attacks. This highlights DROP’s ability to eliminate embedded backdoors that persist in the model long after the attacker has left, outperforming other defenses that fail to mitigate these advanced threats.
V-C Robustness - Accuracy Tradeoff
@@
While DROP demonstrates strong effectiveness against backdoor attacks, it comes with a notable trade-off in MTA compared to some baselines. For example, in Table III with a DPR of 1.25%, DROP achieves an MTA of 74.74%, which is lower than FLTrust (87.93%) and Median (87.14%). Similarly, in Table VI with an MCR of 20%, DROP records an MTA of 68.74%, compared to 86.88% for FLTrust. This reduction in accuracy primarily stems from the strict distillation-based mechanism, which aims to ensure suppression of adversarial influences. Notably, unlike clustering or aggregation-only defenses, DROP uses model stealing (i.e., MAZE) to apply knowledge distillation, which contributes significantly to the observed MTA drop while exhibiting strong robustness. The trade-off between robustness and accuracy is a well-documented challenge in adversarial machine learning [57].
@@
DROPlet - Lightweight Variant of DROP. DROPlet (Section IV-D) is designed for environments where computational efficiency is a priority. Unlike DROP, DROPlet omits the knowledge distillation component, thereby eliminating the associated computational overhead. This makes it particularly suitable for FL deployments with limited resources or scenarios where minimizing server-side processing time is critical. Notably, execution time for a single non-IID run of DROPlet is comparable to other baselines, completing in approximately 30 minutes, whereas DROP takes 2 hours. This minimal increase in runtime comes with the added benefit of robust protection against attacks, whereas some defenses like FLIP take up to 57 hours for similar setups.
While DROPlet may not perform as well as DROP against certain highly stealthy, low-DPR attacks (e.g., CIFAR-10, 1.25% DPR in Table III) or high-MCR attacks (e.g., CINIC-10, 40% MCR in Table VI), it still provides strong protection in many practical configurations, achieving near-zero ASR in less covert attack scenarios and maintaining robustness across a broad range of learning setups (Tables III and VI).
While DROPlet provides strong protection in many scenarios, we acknowledge that an effective defense should remain robust regardless of the adversarial strategy. This limitation becomes especially evident in extremely stealthy attacks such as Chameleon (e.g., CIFAR-10, 2.5% DPR in Table IV), where DROPlet’s effectiveness drops significantly. Nevertheless, its lightweight design offers significant practical value as a plugin to existing FL defenses, requiring minimal code modifications while enhancing robustness without incurring substantial overhead. By striking a balance between computational efficiency and robustness, DROPlet provides a viable defense solution for resource-constrained FL settings.
@@
TABLE VII: MTA and ASR across rounds for various defenses for non-IID data with varying $Dir(\alpha)$, for 1.25% DPR and 20% MCR, using configuration C4 for CIFAR-10 and C9 for EMNIST. FLAME fluctuates wildly and often ends up with models that have not learned anything effectively, leading to random-guess MTA (red) and consequently, zero ASR. Results for FLIP are shown only for $Dir(\alpha)=1$ due to its extreme computational overhead, as the method requires adversarial training for each client at every round. As observed, DROP is the most resilient defense on EMNIST and all defenses struggle on the more complex CIFAR-10 learning task.
V-D Hardness Under Non-IID Settings
@@
Table VII presents the results of evaluating various defenses under increasingly non-IID conditions for CIFAR-10 and EMNIST. While both datasets are commonly used in FL research, CIFAR-10 poses a greater challenge due to its higher intra-class diversity, which allows malicious updates to blend more seamlessly with legitimate ones, making it harder to detect and defend against backdoor attacks. This complexity is reflected in the relatively poor performance of most defenses on CIFAR-10 under non-IID conditions. FLIP for instance, while reducing ASR compared to most other defenses, still suffers from significant fluctuations in ASR, sometimes exceeding 60% under non-IID conditions. This variability highlights the challenge of achieving reliable defense in high-stakes applications where consistency is crucial.
@@
While our defense performs well in many non-IID scenarios, the fluctuating ASR observed in more complex tasks like CIFAR-10 underscores the general difficulty of defending against backdoor attacks under such conditions. These results suggest that more research is needed to develop defenses capable of better identifying malicious updates in the presence of significant data heterogeneity.
VI Conclusion
@@
To address this gap, we propose DROP, a novel defense mechanism that is agnostic to learning configurations and the adversary’s stealthiness. By using this novel architecture based on clustering, activity tracking, and knowledge distillation, DROP offers a comprehensive solution for mitigating targeted backdoor attacks, ensuring reliable protection across diverse FL setups. Our extensive evaluation across multiple datasets, varying data distributions, attacks, and a wide range of attacker configurations highlights DROP’s robustness. Unlike existing baselines, which often fail under certain learning configurations or in the presence of stealthy attacks, DROP consistently achieves near-zero attack success rate.